# PDP: Physics-Based Character Animation via Diffusion Policy

### v1

@@
Abstract
Generating diverse and realistic human motion that can physically interact with an environment remains a challenging research area in character animation. Meanwhile, diffusion-based methods, as proposed by the robotics community, have demonstrated the ability to capture highly diverse and multi-modal skills. However, naively training a diffusion policy often results in unstable motions for high-frequency, under-actuated control tasks like bipedal locomotion due to rapidly accumulating compounding errors, pushing the agent away from optimal training trajectories. The key idea lies in using RL policies not just for providing optimal trajectories but for providing corrective actions in sub-optimal states, giving the policy a chance to correct for errors caused by environmental stimulus, model errors, or numerical errors in simulation. Our method, Physics-Based Character Animation via Diffusion Policy (PDP), combines reinforcement learning (RL) and behavior cloning (BC) to create a robust diffusion policy for physics-based character animation. We demonstrate PDP on perturbation recovery, universal motion tracking, and physics-based text-to-motion synthesis.
Index Terms:
@@ Abstract
Figure 1: PDP performs well across a diverse range of physics-based application domains. Top: Motion tracking. PDP is capable of tracking difficult and highly dynamic motions such as handstands and cartwheels. Middle: Text-to-motion. PDP is also capable of following user-provided text instructions. Bottom: Robustness to perturbations. PDP learns robust recovery strategies from random perturbations.
I Introduction
@@
Developing a framework capable of generating diverse human movements that can proficiently traverse and interact with the environment is a crucial objective in character animation with broad applications in robotics, exoskeletons, virtual/augmented reality, and video games. Many of these applications demand not only a diverse range of human kinematic poses, but also the physical actions needed to achieve them. Previous works have demonstrated that physics-based control tasks can be formulated as a Markov Decision Process and solved through a Reinforcement Learning (RL) algorithm, or as a regression problem and solved using supervised learning techniques such as Behavior Cloning (BC). Despite the achievements observed in dynamic motor skills learning through both RL and BC methodologies, they encounter difficulties in effectively capturing the diversity and multi-model characteristics inherent in human motions.
@@
To address this issue, recent works have explored various generative models. While Generative Adversarial Network (GAN) and Conditional Variational Autoencoder (C-VAE) based methods have been used to capture humanoid skills, GAN based methods can only cover a small set of skills due to mode collapse and VAE-based models suffer from sensitive trade-off between diversity and robustness. Although diffusion models have been used to generate varied kinematic human motion [1, 2], their application to robotics is more recent. Recent robotics research shows that BC combined with diffusion models can effectively learn diverse and multi-modal actions for real-world execution [3, 4]. However, naively training a diffusion-based BC policy is ineffective for physics-based character animation due to compounding errors in high-frequency, under-actuated control tasks, exacerbating the domain shift problem. This issue is especially prominent in bipedal locomotion, where accumulated errors can quickly lead to falling. Can we combine the strengths of RL and diffusion-based BC policy, such that a physically simulated character can perform a diverse set of tasks robustly against distribution shifts due to disturbances in the environment, error in model prediction, or numerical errors in simulation?
@@
We introduce PDP, a novel method that learns a robust diffusion policy for physics-based character animation, addressing the noted challenges. PDP leverages diffusion policies [5] and large-scale motion datasets to learn diverse and multimodal motor skills through supervised learning and diffusion models. To overcome sensitivity to domain shifts, PDP uses expert RL policies to gather physically valid sequences of observations and actions. However, using RL for data collection alone does not resolve domain shift sensitivity.
@@
Our key insight is that RL policies provide not only optimal trajectories but more importantly corrective actions from sub-optimal states. We employ a sampling strategy from robotics literature [6], collecting noisy-state clean-action paired trajectories to train the diffusion policy. We find that the choice of pairing noisy state with clean actions is a critical detail that contributes to producing a robust policy, outperforming the standard clean-state-clean-action trajectory collection and noisy-state-noisy-action sampling strategies for domain randomization. Additionally, we can now pool together data collected by small-task RL policies which can be efficiently trained, and leave the learning of diverse tasks on large-scale datasets to supervised learning.
@@
PDP is a versatile method applicable to various motion synthesis tasks and agnostic to training datasets. We evaluate PDP on locomotion control under large physical perturbations, universal motion tracking, and physics-based text-to-motion synthesis, using different motion capture datasets for each application. Our model captures the multi-modality of human push recovery behavior, outperforming VAE-based methods and deterministic multi-layer perceptron networks on the bump-em dataset [7]. It can also track $98.9\%$ of all AMASS motions and [8] generate motion from textual descriptions [9]. Our contributions are as follows:
@@
II Related Work
II-A Physics-Based Character Animation via RL
@@
In physics-based character animation, the central challenge is developing systems that can learn diverse movements and motor skills for specific tasks. This includes motion tracking, motion generation, and task-oriented applications, which require consideration of physical interactions. Proficiency in these skills is crucial for applications in gaming, virtual reality, and robotics, where authentic and adaptable movements play a vital role.
Single motion imitation using RL enables policies to imitate specific movements or skills, such as walking, running, or jumping [10]. However, these policies alone may not suffice to complete tasks that involve multiple skills. By conditioning policies on future motions, multiple skills can be learned, and a motion generator can guide the policies to complete desired tasks [11, 12]. However, learning control policies that scale to many skills is challenging. [13] proposes to use a mixture of experts, where different experts specialize in different skills. [14] proposes an iterative approach to learning a large number of skills sequentially.
@@
Another challenge in physics-based character animation is capturing the diversity of skills in the motion data. Human behavior is multimodal, meaning a range of plausible behaviors can be employed in the same situation, which can lead to mode collapse, where only a single behavior is learned [15]. To capture the diversity of human skills, Conditional Variational Autoencoders (C-VAE) [16, 17, 18, 19] are used to learn a latent space of skills, enabling random sampling to induce a wide range of motions. [20] achieved a similar capability by combining a diversity reward and adversarial reward that encourage the policy to capture the distribution of human motion data.
Achieving the capability to scale to a large motion dataset while maintaining diversity is still a challenging problem. We propose to use the diffusion model to handle this challenge and demonstrate control policies that perform a large number of skills while maintaining diversity.
II-B Diffusion Models for Motion Synthesis and Robotics
@@
Similar to VAEs, Diffusion models represent another category of generative AI and have exhibited success in the domain of kinematic motion synthesis, showcasing the capability of generating diverse and intricate human motion patterns [1, 21]. Recently, Diffusion Policy [3] has been applied diffusion models to robotic manipulation tasks, human-robot collaborative endeavors [22], and tasks involving following language instructions [23]. These models have primarily concentrated on high-level motion planning with a limited action space, such as forecasting the end-effector trajectory. While effective in stable environments, the application of Diffusion Policy to scenarios where minor inaccuracies in model predictions could result in failure, as observed in physics-based character animation, has only recently been explored in [4]. In contexts requiring high-frequency control, ensuring that the policy remains unaffected by domain shifts is crucial. Through our sampling technique, we have enhanced the robustness of the policy, addressing the domain shift issue and enabling the utilization of diffusion models in the character animation domain. Similar observations have been reported in concurrent studies in quadruped and bi-pedal robot domains [4].
III Methods
@@
Our method consists of three stages. First, we train a set of expert policies, each specialized in a small task but together completing a wide variety of motion tracking tasks in a physics simulator. Second, we generate state-action trajectories from the trained policies stochastically to build a dataset with noisy-state and clean-action trajectories. Lastly, we train a diffusion model via Behavior Cloning (BC) to obtain a single policy that can perform all tasks. Fig. 2 depicts an overview of our method.
III-A Expert Policy Training
@@
III-B Stochastic Data Collection
In the second stage, we utilize the expert policies to generate a dataset for BC. For each task $\mathcal{T}_{i}$, we create a dataset $\mathcal{D}_{\mathcal{T}_{i}}$ by rolling out policy $\pi_{\mathcal{T}_{i}}$ and collecting trajectories. Specifically, we sample a motion task $\tau\in\mathcal{T}_{i}$, and run the policy to generate a sequence $\{\bm{o}_{0},\bm{a}_{0},\bm{o}_{1},\bm{a}_{1},\dots,\bm{o}_{N},\bm{a}_{N}\}$, where $\bm{a}_{t}=\pi_{\mathcal{T}_{i}}(\bm{o}_{t},\tau)+\epsilon$ is a noisy version of optimal action proposed by the expert policy. The tuples $(\bm{o}_{t},\tau,\pi_{\mathcal{T}_{i}}(\bm{o}_{t},\tau))$ which correspond to the observation, task/goal information, and action are added to the dataset $\mathcal{D}_{\mathcal{T}_{i}}$. We repeat the data collection process until a maximum number of data points are collected, and use $\mathcal{D}=\bigcup\mathcal{D}_{\mathcal{T}_{i}}$ as the dataset for BC. Note that the optimal action, not the noisy action $a_{t}$, is stored in $\mathcal{D}_{\mathcal{T}_{i}}$ and can be thought of as a corrective action from a noisy observation. This important detail, inspired by the DASS strategy proposed by [6], results in a training set that consists of sequences of noisy-state and clean-action pairs. This allows the collected data to cover a wider range of observation space compared to naively collecting clean optimal state-action trajectories, effectively creating a “noise band” around the clean trajectories. Another potential option is to collect noisy-state and noisy-action pairs, a common domain randomization practice in robotics to battle the sim-to-real gap. We found this randomization strategy produces less robust policies for character animation.
Figure 2: PDP Overview. Top: First, we train expert RL policies $\mathcal{\pi}_{\mathcal{T}_{i}}$ on tasks $\mathcal{T}_{i}$. We use $\mathcal{\pi}_{\mathcal{T}_{i}}$ to create a dataset of noisy-state clean-actions. We then use BC to train a diffusion model. Bottom: Our model is a transformer encoder-decoder architecture. Block-B is used for text conditioned tasks, while all other tasks use Block A.
III-C Behavior Cloning with Diffusion Policy
@@
III-C1 Diffusion model
We employ Diffusion Policy [3], which models the action distribution conditioned on the observations as a denoising process using a Denoising Diffusion Probabilistic Model (DDPM) [24]. Given a dataset of sequences $\mathcal{D}$ collected using the method described previously, the denoising process is learned by a noise-prediction network $\epsilon_{\theta}(\bm{A}_{t}^{k},\bm{O}_{t},\bm{\tau}_{t},k)$, where $\bm{A}_{t}^{k}$ is an action sequence sampled from $\mathcal{D}$ with added Gaussian noise and $k$ is the diffusion step. The diffusion model is conditioned on $\bm{O}_{t}$, the corresponding observation sequence. $\tau$ is the necessary task or goal information, and $\theta$ is the set of learned model parameters. $\bm{A}_{t}^{k}$ and $\bm{O}_{t}$ are length $T$ sequences of actions and observations beginning at episode timestep $t$. Sampling is then achieved through a denoising process known as Stochastic Langevin Dynamics [25] starting from pure random noise.
$\bm{A}_{t}^{k-1}=\alpha(\bm{A}_{t}^{k}-\gamma\epsilon_{\theta}(\bm{A}_{t}^{k},\bm{O}_{t},\bm{\tau}_{t},k)+\mathcal{N}(0,\sigma^{2}\bm{I})),$ (1)
@@ III-C1 Diffusion model
$\mathcal{L}=MSE(\epsilon^{k},\epsilon_{\theta}(\bm{A}_{t}^{0}+\epsilon^{k},\bm{O}_{t},\bm{\tau}_{t},k)),$ (2)
@@ III-C2 Model Architecture
Fig. 2 depicts our model architecture. We adopt a similar architecture to the time-series diffusion transformer proposed by Diffusion Policy [3], with slight modifications depending on the application. For locomotion control and motion tracking, task information is included in the observation. For text-to-motion, the conditioning for the action sequence is computed as follows: a raw text prompt is encoded using the CLIP ViT-B/32 model [26] , then passed through an MLP text encoder. The observation $\bm{O}{t}$ is also fed through an MLP encoder. The diffusion step $k$ is embedded into the same space and added to the text embedding. This result is fed through a Feature-wise Linear Modulation (FiLM) layer [27], which applies a learned element-wise scale and shift transformation to the embedding of $\bm{O}{t}$. Finally, the diffusion step embedding is concatenated with the FiLM layer result to produce our condition, serving as the input to the transformer encoder. This conditioning method is represented in Block B of Figure 2. The transformer decoder then takes an embedding of the noisy action sequence $\bm{A}_{t}^{k}$ along with the encoder result and predicts the noise applied to the action $\epsilon^{k}$. For motion tracking tasks, we condition our transformer encoder using Block A of Figure 2, which is similar to Block B but without the CLIP text.
IV Experimental Details

### v2

@@
Abstract.
Generating diverse and realistic human motion that can physically interact with an environment remains a challenging research area in character animation. Meanwhile, diffusion-based methods, as proposed by the robotics community, have demonstrated the ability to capture highly diverse and multi-modal skills. However, naively training a diffusion policy often results in unstable motions for high-frequency, under-actuated control tasks like bipedal locomotion due to rapidly accumulating compounding errors, pushing the agent away from optimal training trajectories. The key idea lies in using RL policies not just for providing optimal trajectories but for providing corrective actions in sub-optimal states which gives the policy a chance to correct for errors caused by environmental stimulus, model errors, or numerical errors in simulation. Our method, Physics-Based Character Animation via Diffusion Policy (PDP), combines reinforcement learning (RL) and behavior cloning (BC) to create a robust diffusion policy for physics-based character animation. We demonstrate PDP on perturbation recovery, universal motion tracking, and physics-based text-to-motion synthesis.
@@ Abstract.
submissionid: 1034copyright: rightsretainedconference: SIGGRAPH Asia Conference Papers 2024; December 03–06, 2024; Tokyo, Japanjournalyear: 2024publicationmonth: 12isbn: 979-8-4007-1131-2s/24/12doi: 10.1145/3680528.3687683article: 86journal: TOGccs: Computing methodologies Animationccs: Computing methodologies Physical simulation
@@
Figure 1. PDP performs well across a diverse range of physics-based application domains. Top: Motion tracking. PDP is capable of tracking difficult and highly dynamic motions such as handstands and cartwheels. Middle: Text-to-motion. PDP is also capable of following user-provided text instructions. Bottom: Robustness to perturbations. PDP learns robust recovery strategies from random perturbations.
1. Introduction
@@
Developing a framework capable of generating diverse human movements that can traverse and interact with the environment is a crucial objective in character animation with broad applications in robotics, exoskeletons, virtual/augmented reality, and video games. Many of these applications demand not only a diverse range of human kinematic poses, but also the physical actions needed to achieve them. Previous works have demonstrated that physics-based control tasks can be formulated as a Markov Decision Process and solved through a Reinforcement Learning (RL) algorithm, or as a regression problem and solved using supervised learning techniques such as Behavior Cloning (BC). Despite the achievements observed in dynamic motor skills learning through both RL and BC methodologies, they encounter difficulties in effectively capturing the diversity and multi-model characteristics inherent in human motions.
@@
To address this issue, recent works have explored various generative models. While Conditional Variational Autoencoders (C-VAEs) and Generative Adversarial Networks (GANs) have been used to capture humanoid skills, VAE-based models suffer from sensitive trade-off between diversity and robustness, while GAN-based methods often suffer from mode collapse without additional objectives (Dou et al., 2023). Although diffusion models have been used to generate varied kinematic human motion (Tevet et al., 2023; Tseng et al., 2022), their application in high-frequency control domains is relatively unexplored. Recent work in robotics shows that BC combined with diffusion models can effectively learn diverse and multi-modal actions for real-world execution (Chi et al., 2023; Huang et al., 2024). However, naively training a diffusion-based BC policy is ineffective for physics-based character animation due to compounding errors in high-frequency, under-actuated control tasks, exacerbating the domain shift problem. This issue is especially prominent in bipedal locomotion, where accumulated errors can quickly lead to falling. Can we combine the strengths of RL and diffusion-based BC policy, such that a physically simulated character can perform a diverse set of tasks robustly against distribution shifts due to disturbances in the environment, error in model prediction, or numerical errors in simulation?
@@
We introduce PDP, a novel method that learns a robust diffusion policy for physics-based character animation, addressing the noted challenges. PDP leverages diffusion policies (Song and Ermon, 2020) and large-scale motion datasets to learn diverse and multimodal motor skills through supervised learning and diffusion models. To overcome sensitivity to domain shifts, PDP uses expert RL policies to gather physically valid sequences of observations and actions. However, using RL for data collection alone does not resolve domain shift sensitivity.
@@
Our key insight is that RL policies provide not only optimal trajectories but more importantly corrective actions from sub-optimal states. We employ a sampling strategy from robotics literature (Xie et al., 2020), collecting noisy-state clean-action paired trajectories to train the diffusion policy. We find that the choice of pairing noisy state with clean actions is a critical detail that contributes to producing a robust policy, outperforming the standard clean-state-clean-action trajectory collection and noisy-state-noisy-action sampling strategies for domain randomization. Additionally, we can now pool together data collected by small-task RL policies which can be efficiently trained, and leave the learning of diverse tasks on large-scale datasets to supervised learning.
@@
PDP is a versatile method applicable to various motion synthesis tasks and agnostic to training datasets. We evaluate PDP on locomotion control under large physical perturbations, universal motion tracking, and physics-based text-to-motion synthesis, using different motion capture datasets for each application. Our model captures the multi-modality of human push recovery behavior, outperforming VAE-based methods and deterministic multi-layer perceptron networks on the Bump’em dataset (Werling et al., 2023). It can also track $98.9\%$ of all AMASS motions and (Mahmood et al., 2019) generate motion from textual descriptions (Guo et al., 2022). Our contributions are as follows:
@@
2. Related Work
2.1. Physics-Based Character Animation
@@
In physics-based character animation, the central challenge is developing systems that learn a diverse range of realistic motions. Such motions manifest in motion tracking, motion generation, and task-oriented applications that require consideration of physical interactions.
Methods for learning individual or relatively similar motions, such as walking, running, or jumping are well-established (Peng et al., 2018, 2021). However, these methods often do not suffice to complete tasks that involve multiple skills. Other methods have been successful in learning motion tracking policies capable of multiple skills using model-free RL (Bergamin et al., 2019; Park et al., 2019) and model-based RL (Fussell et al., 2021). However, scaling to large diverse datasets is challenging. UniCon (Wang et al., 2020) introduced a motion tracking controller that scales to large and diverse motion datasets by using a novel constrained multi-objective reward function. Other works propose a mixture of experts (Won et al., 2020a), where different experts specialize in different skills. PHC (Luo et al., 2023) proposes an iterative approach to learning a large number of skills sequentially.
@@
Another challenge in physics-based character animation is capturing diversity in motion data for use in downstream tasks. Human behaviors are multimodal, meaning a range of plausible behaviors can be employed in the same situation. A common method of capturing diversity in motions is to employ a Variational Autoencoder (VAE) to learn a latent space of skills, then sampling from the VAE prior to produce a wide range of motions (Yao et al., 2022, 2023; Zhu et al., 2023; Won et al., 2022; Merel et al., 2018). These latent motion representations can then be used for downstream tasks such as motion generation (Luo et al., 2024) or object iterations (Merel et al., 2020). Adversarial methods have also been proposed for capturing motion diversity (Peng et al., 2022; Dou et al., 2023), which combine a diversity reward and adversarial reward that encourage the policy to mimic the motion distribution.
Robustness issues also arise in behavior cloning methods where error accumulation can easily push the policy out of distribution. One method for improving policy robustness is to continually roll out the current policy, collecting on-policy data to train a student in the next learning iteration, as in DAgger (Ross et al., 2011). Alternatively, robustness can be achieved by injecting perturbations into the state-action pairs in the training dataset, effectively expanding the distribution of states seen during training, similar to DASS (Xie et al., 2020).
2.2. Diffusion Models for Motion Synthesis and Robotics
@@
Similar to VAEs, Diffusion models represent another category of generative AI and have exhibited success in the domain of kinematic motion synthesis, showcasing the capability of generating diverse and intricate human motion patterns (Tevet et al., 2023; Tseng et al., 2023). Recently, Diffusion Policy (Chi et al., 2023) has effectively applied diffusion models to robotic manipulation tasks, human-robot collaborative endeavors (Ng et al., 2024), and tasks involving following language instructions (Zhang et al., 2022). These models have primarily concentrated on high-level motion planning with a limited action space, such as forecasting the end-effector trajectory. While effective in low-frequency environments, the application of Diffusion Policy to high-frequency scenarios where minor inaccuracies in model predictions could result in failure, such as in physics-based character animation, remains relatively unexplored. Concurrent work, DiffuseLoco (Huang et al., 2024), is similar to PDP, employing a diffusion model to distill an offline dataset of multimodal skills, however they focus on simple locomotion gaits due to their policy being deployed on a real robot.
3. Methods
@@
Our method consists of three stages. First, we train a set of expert policies, each specialized in a small task but together completing a wide variety of motion tracking tasks in a physics simulator. Second, we generate state-action trajectories from the trained policies stochastically to build a dataset with noisy-state and clean-action trajectories. Lastly, we train a diffusion model via Behavior Cloning (BC) to obtain a single policy that can perform all tasks. Fig. 2 gives an overview of our system.
3.1. Expert Policy Training
@@
3.2. Stochastic Data Collection
In the second stage, we utilize the expert policies to generate a dataset for BC. For each task $\mathcal{T}_{i}$, we create a dataset $\mathcal{D}_{\mathcal{T}_{i}}$ by rolling out policy $\pi_{\mathcal{T}_{i}}$ and collecting trajectories. Specifically, we sample a motion task $\tau\in\mathcal{T}_{i}$, and run the policy to generate a sequence $\{\bm{o}_{0},\bm{a}_{0},\bm{o}_{1},\bm{a}_{1},\dots,\bm{o}_{N},\bm{a}_{N}\}$, where $\bm{a}_{t}=\pi_{\mathcal{T}_{i}}(\bm{o}_{t},\tau)+\epsilon$ is a noisy version of optimal action proposed by the expert policy. The tuples $(\bm{o}_{t},\tau,\pi_{\mathcal{T}_{i}}(\bm{o}_{t},\tau))$ which correspond to the observation, task/goal information, and action are added to the dataset $\mathcal{D}_{\mathcal{T}_{i}}$. We repeat the data collection process until a maximum number of data points are collected, and use $\mathcal{D}=\bigcup\mathcal{D}_{\mathcal{T}_{i}}$ as the dataset for BC. Note that the optimal action, not the noisy action $a_{t}$, is stored in $\mathcal{D}_{\mathcal{T}_{i}}$ and can be thought of as a corrective action from a noisy observation. This important detail, inspired by the DASS strategy proposed by (Xie et al., 2020), results in a training set that consists of sequences of noisy-state and clean-action pairs. This allows the collected data to cover a wider range of observation space compared to naively collecting clean optimal state-action trajectories, effectively creating a ”noise band” around the clean trajectories. Our method extends DASS by further widening the noise band. Specifically, we generate short recovery episodes by initializing the character with a random root position and orientation offset from its original motion and allow it to recover to the original motion over several timesteps. This approach applies the noise band not only to the joints but also to the character’s overall pose, helping to mitigate drift over time.
Another potential option for sampling is to collect noisy-state and noisy-action pairs, a common domain randomization practice in robotics to battle the sim-to-real gap. We found this randomization strategy produces less robust policies for character animation.
Figure 2. PDP Overview. Top: First, we train expert RL policies $\mathcal{\pi}_{\mathcal{T}_{i}}$ on tasks $\mathcal{T}_{i}$. We use $\mathcal{\pi}_{\mathcal{T}_{i}}$ to create a dataset of noisy-state clean-actions. We then use BC to train a diffusion model. Bottom: Our model is a transformer encoder-decoder architecture. Block-B is used for text-conditioned applications, while other applications use Block A. Note that these applications are trained separately on their own distilled dataset.
3.3. Behavior Cloning with Diffusion Policy
@@
3.3.1. Diffusion model
We employ Diffusion Policy (Chi et al., 2023), which models the action distribution conditioned on the observations as a denoising process using a Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020). Given a dataset of sequences $\mathcal{D}$ collected using the method described previously, the denoising process is learned by a noise-prediction network $\epsilon_{\theta}(\bm{A}_{t}^{k},\bm{O}_{t},\bm{\tau}_{t},k)$, where $\bm{A}_{t}^{k}$ is an action sequence sampled from $\mathcal{D}$ with added Gaussian noise and $k$ is the diffusion step. The diffusion model is conditioned on $\bm{O}_{t}$, the corresponding observation sequence. $\tau$ is the necessary task or goal information, and $\theta$ is the set of learned model parameters. As in Diffusion Policy, $\bm{A}_{t}^{k}$ is a length $T$ sequence of actions beginning at timestep $t$ and $\bm{O}_{t}$ is a length $T$ sequences of observations ending at timestep $t$. Sampling is then achieved through a denoising process known as Stochastic Langevin Dynamics (Welling and Teh, 2011) starting from pure random noise.
(1) $\bm{A}_{t}^{k-1}=\alpha(\bm{A}_{t}^{k}-\gamma\epsilon_{\theta}(\bm{A}_{t}^{k},\bm{O}_{t},\bm{\tau}_{t},k)+\mathcal{N}(0,\sigma^{2}\bm{I})),$
@@ 3.3.1. Diffusion model
(2) $\mathcal{L}=MSE(\epsilon^{k},\epsilon_{\theta}(\bm{A}_{t}^{0}+\epsilon^{k},\bm{O}_{t},\bm{\tau}_{t},k)),$
@@ 3.3.2. Model Architecture
Figure 2 depicts our model architecture. We adopt a similar architecture to the time-series diffusion transformer proposed by Diffusion Policy (Chi et al., 2023), with slight modifications depending on the application. For locomotion control and motion tracking, task information is included in the observation. For text-to-motion, the conditioning for the action sequence is computed as follows: a raw text prompt is encoded using the CLIP ViT-B/32 model (Radford et al., 2021) , then passed through an MLP text encoder. The observation $\bm{O}{t}$ is also fed through an MLP encoder. The diffusion step $k$ is embedded into the same space and added to the text embedding. This result is fed through a Feature-wise Linear Modulation (FiLM) layer (Perez et al., 2017), which applies a learned element-wise scale and shift transformation to the embedding of $\bm{O}{t}$. Finally, the diffusion step embedding is concatenated with the FiLM layer result to produce our condition, serving as the input to the transformer encoder. This conditioning method is represented in Block B of Figure 2. The transformer decoder then takes an embedding of the noisy action sequence $\bm{A}_{t}^{k}$ along with the encoder result and predicts the noise applied to the action $\epsilon^{k}$. For motion tracking tasks, we condition our transformer encoder using Block A of Figure 2, which is similar to Block B but without the CLIP text.
4. Experiments